A man has a 40 watt and two 60 watt bulbs in a room. How much will it cost him to keep them lit for 8 hours per day in the month of June if the cost of a unit is #20.
1
Expert's answer
2020-06-29T14:14:24-0400

Assuming than unit is 1kWh (kilowatt-hour).

Cost for light:

"Cost=Q\times UnitPrice"

where Q is quantity of energy spent.

"Q=P\times t"

where P is power of all bulbs (in kilowatts), t - time which bulbs were used (in hours).

"t=t_{1}\times N_{d}"

where t1 - hours during 1 day which bulbs were used, Nd - number of days in June.

So we have

"Cost=P\times t_{1}\times N_{d} \times UnitPrice"

"Cost=(0.040+0.060+0.060)\times 8 \times 30 \times 20 =" 768